# Rapid near-inertial wave generation at topography in response to surface wind forcing

### v2

@@ 3 Numerical model and analysis methodology
For our control experiment, a Coriolis parameter of $f=-10^{-4}$ s-1 corresponding to the Southern mid-latitudes is selected, and a 1000 m tall Gaussian ridge with full width at half maximum (FWHM) $W=10~{}\mathrm{km}$, running north to south through the domain, is placed in the centre. These parameters are chosen to be consistent with a large abyssal hill, such as those found in the Southern Ocean (Goff and Arbic, 2010). The isopycnal model has 21 layers whose thicknesses add up to a total depth $H=4000\;\mathrm{m}$. The top layer mimics a mixed layer and has a thickness of $50~{}\mathrm{m}$ in the default case; the second layer below mimics a thermocline and initially has a thickness of $150~{}\mathrm{m}$; the remaining 19 layers below have an initial thickness of $200~{}\mathrm{m}$. The top layer has a density of $1018\;\mathrm{kg}/\mathrm{m}^{3}$, followed by large jump in density with an equivalent buoyancy frequency $N=200|f|$ to simulate a thermocline. The remaining 19 layers have a linearly increasing density to give a constant buoyancy frequency of $40|f|$, an approximation of a typical buoyancy frequency observed beneath the thermocline (Reagan, James R. et al., 2021).
@@
where $\tau_{0}$ is the peak wind stress, $L_{\tau}=2\sqrt{2\log{(2)}}\sigma$ is the FWHM of the wind stress, and $T_{\tau}$ the forcing duration, with default values of $T_{\tau}=10\;\mathrm{hours}$ and $L_{\tau}=80\;\mathrm{km}$, consistent with observed tropical cyclone size distributions (Zhang and Chan, 2023). The default wind stress value $\tau_{0}=10\;\mathrm{N}\,\mathrm{m}^{-2}$ corresponds to the strength of a category 4 tropical cyclone with surface wind speed of $63\;\mathrm{m}\,\mathrm{s}^{-1}$. The wind pulse is focused at the meridional-centre of the domain, and the eastward and westward oscillation serves to force the surface layer without inducing a significant time-mean flow, providing a robust idealised model of the process of interest (i.e., high frequency wind forcing by storms). A small time-mean flow of less than $10^{-3}\;\mathrm{mm}\,\mathrm{s}^{-1}$ to the west is still present, but is too small to meet the $|f|<kU_{0}$ condition required to support topographic lee waves.
@@
Employing this symmetry, we calculate the wave energy fluxes propagating outwards through a square centred on $x_{0}$, $y_{0}$ with side lengths $D$. The surface generated NIW energy flux, $E_{S}$, corresponds to the energy flux propagating meridionally through the northern and southern sides, whereas the NIWs resulting from interactions with topography, $E_{T}$, correspond to the wave energy fluxes through the eastern and western boundaries of the square. Energy fluxes are the product of the velocities normal to the face, $u$ or $v$, and pressure $P$. The symmetry of the domain ensures that the energy of the topographically generated near-inertial waves that are the focus of this study is captured within $E_{T}$.
@@
In order to isolate the velocities $u$ and pressures $P$ required to calculate $E_{T}$, we subtract reference values from $2000\mathrm{km}$ west of the ridge, such that $u^{\prime}=u-u_{\rm far}$ and likewise for $P$. Multiplying the perturbation velocity and pressure, and summing over the opposing sides of the square yields the zonal $E_{i,T}$ and meridional $E_{i,S}$ energy fluxes at the $i$th layer:
$E_{i,T}(y,t)$ $=\hskip 14.22636pt\left[P_{i}^{\prime}u_{i}^{\prime}\right]_{x=\frac{D}{2}}\hskip 14.22636pt-\hskip 14.22636pt\left[P_{i}^{\prime}u_{i}^{\prime}\right]_{x=-\frac{D}{2}},$ (7)
[…]
These energy fluxes are then averaged in space and time, to determine the mean quantities $\overline{E_{T}}$ and $\overline{E_{S}}$:
@@ 4 Results
We first provide a qualitative demonstration of the proposed wind-driven topographic near-inertial waves. Figure 3 shows the meridional velocities and isopycnal height anomalies associated with the meridionally propagating near-inertial waves generated directly at the surface. Internal waves are immediately visible through the full water column, similar to the analytic theory shown in Fig. 1. As time progresses, we see that the mode 1 waves (characterised by the single zero crossing) dominant in panel (a) give way to third and fourth vertical modes in panels (c) and (d). This behaviour is consistent with the modal de-phasing theory in that the faster, lower modes appear first, which propagate more rapidly from the centre of the storm. By inspecting the phase shifts of individual vertical modes and the beam-like structures, it is clear that the frequency is near-inertial. Importantly to the generation of waves from topography, Fig. 3 shows significant motions at the seafloor immediately after the storm event.
To visualise the wind-driven topographic near-inertial waves, the velocity and interfacial height anomalies along two zonal transects are plotted in Fig. 4. Here, as in Eq. (7), the anomalies are calculated with respect to their far-field values to remove the effects of non-topographically-generated waves. In all panels, clear beams can be seen propagating upwards and outwards from the top of the ridge. The two rows, corresponding to locations $150$ and $200~{}\mathrm{km}$ northwards from the centre of the wind forcing, show meridional variations to the waves structure and phase. This is due to the time taken for the de-phasing signal to propagate outwards from the storm, as well as the meridional structure of the storm itself. Similarly to Fig. 3, the near-inertial nature of these waves is clear from the phase reversals every half inertial period.
@@
The near-inertial nature of the bottom velocity is more clearly visible in Fig. 5a, which shows the bottom velocity directly beneath the wind forcing. Analogous to Fig. 1b, Fig. 5a shows the bottom velocity directly beneath the wind forcing band. Here, the model shows an immediate bottom response. There is a clear super-inertial signal in the bottom flow, which enables the generation of propagating internal waves in the presence of appropriate topography. A Hovmöller diagram in Fig. 5b shows a signal in the resulting zonal velocity propagating away from the ridge, including several superimposed vertical modes, fanning out from the topography.
Figure 5: Both panels show velocities 150 km north of the ridge. Panel a) Bottom zonal velocity 1000 km West of ridge, b) Hovmöller diagram of zonal velocity anomalies 5 layers from the bottom in the vicinity of the ridge. Both panels show velocities 150 km north of the ridge, as in the bottom row of Fig. 4.
The waves visible in Fig. 4 and Fig. 5 have characteristics consistent with the proposed mechanism, and thus serve as a qualitative demonstration of wind-induced topographically generated near-inertial waves. We now quantify the wave energy, in order to assess in what circumstances, if any, these waves might pose a significant source of bottom intensified wave activity.
@@ 4.1 Wave Energy Flux
Having established the existence of wind-induced topographically generated near-inertial waves in our configuration, we now proceed to calculate the energy flux of these waves, as well as for the wind-generated surface NIWs. From (7), the time averaged energy flux in our reference configuration (Figs. 3, 4, and 5 above) associated with the wind-generated topographic NIWs is $0.13\;\mathrm{mW}\,\mathrm{m}^{-2}$, compared to $1.9\;\mathrm{mW}\,\mathrm{m^{-2}}$ for the wind-generated surface NIWs. Thus, the energy flux associated with wind-generated topographic NIWs is 7% the size of that of surface generated NIWs with our default set of parameters.
@@ 4.2 Perturbation experiments
We run experiments varying both the properties of the storm and ridge to assess which parameters the energy flux is most sensitive to and the conditions under which the topographically generated waves are most significant.
Figure 6: The energy flux of the wind-driven topographic near-inertial (zonal) waves when varying a) forcing duration; b) surface forcing wind speed; c) ridge height; d) ridge width. The default values are shown with an $\times$. The energy flux for the wind-driven surface generated near-inertial (meridional) waves is shown in panels a) and b) only, as it is unaffected by varying the shape of the topography.
When altering the duration of the wind forcing pulse (Fig. 6a) there is a peak in both topographic and surface energy fluxes when the forcing duration is less than the inertial period, i.e., $2\pi/|f|\approx 17.5$ hours. This result is consistent with expectations for surface-driven near-inertial waves, as internal wave generation is only supported by super-inertial forcing frequencies. Forcing durations right at the inertial frequency would overlap significantly with the sub-inertial band, whereas a more rapid surface forcing would contain less energy at the highly resonant inertial period.
@@
Finally, Fig. 6d shows the sensitivity of topographic energy flux when varying the ridge width. Here, both the scaled energy flux, as in Eq. (7), is shown alongside the product of the energy flux and width. This plot is designed to show the way that the total wave energy changes, as well as energy flux, which scales inversely to ridge width. Here, we see that the energy flux per unit area (Fig. 6d, left axis) is greatest for the narrowest ridge, but that the total wave energy per metre meridionally (Fig. 6d, right axis) peaks for a $30\mathrm{km}$ wide ridge. This effect is due to the maximum topographic slope becoming shallower than the angle of the wave beam, suppressing the downward propagating part of the wave-field visible in Fig. 4 (result not shown). In Eq. (5), $k$ refers to the dominant wavenumber of the ridge, which in our case is $k\sim 1/W$. This is consistent with the our results in Fig. 6, whereby the energy flux decreases with increasing ridge width. Intuitively, this means that for wider, and in turn less steep topography, the energy flux per unit area of the topography decreases.
@@ 5 Discussion and conclusions
We demonstrate a novel mechanism for the generation of wind-driven near-inertial waves at topography within hours of the passage of a storm. These waves are generated by the rapid de-phasing of horizontally propagating full-depth wave modes induced by a transient wind event, and interaction of the resulting bottom flow with topography to radiate waves. This generation mechanism is showcased using an isopycnal model forced by a wind pulse that mimics a storm. The idealised model setup is carefully constructed to enable a clean separation of our proposed waves from surface generated near-inertial waves and to eliminate any significant time-mean flow that could generate lee waves.
@@
The perturbation experiments shown in Fig. 6 reveal that the variation of wave energy fluxes in our experiment is consistent with the existing internal wave theory. With our control experiment, namely a $1000\;\mathrm{m}$-high, $10\;\mathrm{km}$ wide ridge in a $4000\;\mathrm{m}$ deep ocean, with an $80\;\mathrm{km}$ wide (FWHM) category 4 tropical cyclone overhead, the wind-generated topographic near-inertial waves have energy fluxes of $0.13\;\mathrm{mW}/\mathrm{m}^{2}$, that is, $7\%$ the energy flux in surface generated NIWs in the same experiment. With the most narrow topography that we test, the maximum energy flux attained is $0.39\;\mathrm{mW}/\mathrm{m}^{2}$, or $18\%$ of the surface generated NIWs. However, as such a prominent ridge or seamount would be uncommon, we consider as a more typical range, the energy flux ratios from the strength and duration perturbation experiments (Fig. 6a,b). From these two experiments, the ratios of the topographically generated to the directly surface generated NIW energy fluxes range from 5% at the longest duration and strongest storm, to 10% at the shortest and weakest storms. This range is consistent with the 6% of surface NIW energy estimated to reach the seafloor in the work by Jouanno et al. (2016), except that with our proposed mechanism this energy is communicated rapidly (within 100 hours of the storm) to the bottom.
Wind-driven topographic near-inertial waves may be especially important for abyssal mixing, as they come at near-inertial frequencies and hence exhibit high vertical shear, making them more likely to dissipate close to the ocean bottom than more super-inertial waves like internal tides. Note that in our experiments, wave breaking and turbulence was minimised by the experimental design so that the propagating wave-field was easier to study, and hence we made no attempt to estimate abyssal mixing created by this mechanism. Some characteristics of the proposed wind-driven topographic near-inertial waves — namely bottom generation and propensity to break near the seafloor — are similar to topographic lee waves. Topographic lee waves are well studied, and are associated with between $0.05-0.85~{}\mathrm{mW}/\mathrm{m}^{2}$ of globally averaged energy flux (Bennetts et al., 2024), although they are mostly concentrated in the Southern Ocean. In our study, the energy flux associated with the wind-driven topographic near-inertial waves overlaps this range, even when omitting the most extreme topographic shapes tested. However, the mechanism described in this study relies on specific and short lived circumstances, namely a storm event near prominent topography. Our proposed wave generation mechanism therefore is not expected to be a major contributor to the global internal wave field in a globally averaged sense. Instead, the mechanism could be important for explaining intermittent locally enhanced internal wave generation at the seafloor and deep-ocean mixing.
@@
Recent studies have identified evidence for unexplained enhancement of internal wave activity at the seafloor. van Haren (2020) finds anomalously large bottom near-inertial oscillations, mixing and deep overturning at $3000~{}\mathrm{m}$ depth $1000~{}\mathrm{km}$ away from a category 4 tropical cyclone. This bottom signal’s first detection and intensification happened simultaneously with the approach of the cyclone, pointing to a rapid (on the order of hours or a few days) communication of the storm’s energy to the deep ocean. Similarly, in another observational study, Morozov and Velarde (2008) show a rapid bottom response coinciding with the arrival of two cyclones, in addition to the downward propagating wave-packet which takes two weeks to reach the seafloor (see Figs. 11 and 12 by Morozov and Velarde (2008)). A modelling study of storm-induced flows shows enhanced near-inertial energies at the seafloor within 5 days of the passage of a storm, several days ahead of the downward propagating wave packets (see Fig. 9 by Jouanno et al. (2016)). Jouanno et al. (2016) note that while this bottom NIW generation is linked to the storm, the associated energies are two orders of magnitude smaller than the surface generated NIWs and were thus not investigated further. However, our perturbation experiments suggest this ratio of surface to bottom generated NIW energy flux could be an order of magnitude larger than was found by Jouanno et al. (2016), warranting further investigation. Finally, by reviewing global in-situ observations, Waterhouse et al. (2022) found that large regions of the ocean exhibit excess upgoing-rotary-shear, indicative of bottom NIW activity. This unaccounted for bottom NIW activity lends further support to the need for additional bottom generation pathways for near-inertial waves. We argue that the mechanism proposed in the study could at least partly account for these findings in the existing literature. Further study is needed to clarify the extent to which this may be the case Waterhouse et al. (2022). Here, we studied the generation of wind-driven topographic near-inertial waves in a highly idealised domain to cleanly separated them from other flow features. An important next step is to identify the waves in more realistic simulations. However, delineating topographic near-inertial waves from other flow factors is difficult due to their generation by the de-phasing of the vertical modes of the surface response, which itself includes NIWs of similar frequencies and spatial scales. Our experiments were carefully designed to separate the wind-driven topographic near-inertial waves from other flow features as cleanly as possible. Consequently, this limited the scope of our study to simple dynamics, topographic features, and surface forcing profiles. To investigate topographic near-inertial waves in a more realistic numerical simulation or with observations, these intricacies of disentangling them from other flow features will need to be handled carefully. One option could be to exploit the difference in vertical propagation between the topographic and wind-generated near-inertial waves. If a filter were used to separate the upwards and downwards propagating signals, perhaps following the spectral approach employed by Olbers and Eden (2017) or Waterhouse et al. (2022), one could identify areas of increased vertical energy flux beneath storm systems. However, one would still need to remove the background internal wave fields, among other oscillatory signals that could obscure the waves of interest. To obtain a stronger signal, a statistical approach could be employed, e.g., averaging over many storm events in several locations above regions of varying topographic prominence. At each location, measurements of the bottom mixing could be taken both in the absence and presence of an overhead storm. Future studies could employ this approach on high-resolution model output or observations.

### v4

@@ 3 Numerical model and analysis methodology
The top isopycnal layer mimics a mixed layer and has a thickness of $50~{}\mathrm{m}$ in the default case; the second layer below mimics a thermocline and initially has a thickness of $150~{}\mathrm{m}$; the remaining 19 layers below have an initial thickness of $200~{}\mathrm{m}$. A Coriolis parameter of $f=-10^{-4}$ s-1 corresponding to the Southern mid-latitudes is selected. The top layer has a density of $1018\;\mathrm{kg}/\mathrm{m}^{3}$, followed by a large jump in density with an equivalent buoyancy frequency $N=200|f|$ to simulate a thermocline. The remaining 19 layers have a linearly increasing density to give a constant buoyancy frequency of $40|f|$, an approximation of a typical buoyancy frequency observed beneath the thermocline (Reagan, James R. et al., 2021).
For our control experiment, a zonally symmetric 1000 m tall Gaussian ridge with full width at half maximum (FWHM) $W=10\;\mathrm{km}$, running north to south through the domain, is placed in the domain centre. These parameters for the cross-ridge and vertical scales are chosen to be consistent with a typical seamount (Wessel et al., 2010).
@@
where $\tau_{0}$ is the peak wind stress, $L_{\tau}=2\sqrt{2\log{(2)}}\sigma$ is the FWHM of the wind stress, and $T_{\tau}$ the forcing duration, with default values of $T_{\tau}=10\;\mathrm{hours}$ and $L_{\tau}=80\;\mathrm{km}$, consistent with observed tropical cyclone size distributions (Zhang and Chan, 2023). The default wind stress value $\tau_{0}=10\;\mathrm{N}\,\mathrm{m}^{-2}$ corresponds to the strength of a category 4 tropical cyclone with surface wind speed of $63\;\mathrm{m}\,\mathrm{s}^{-1}$. The wind pulse is focused at the meridional-centre of the domain, and the eastward and westward oscillation serves to force the surface layer without inducing a significant time-mean flow, providing a robust idealised model of the process of interest (i.e., high-frequency wind forcing by storms). A small westward time-mean flow, less than $10^{-6}\;\mathrm{m}\,\mathrm{s}^{-1}$, is still present, but this is too small to meet the $|f|<kU_{0}$ condition required to support topographic lee waves.
In choosing to filter the waves by symmetry, a drawback is that our wind event lacks zonal gradients, restricting the divergence of the wind field to the meridional gradient. This formulation means that, while near-inertial waves are expected to be generated due to this sharp, unidirectional gradient, the magnitude of these waves may be significantly reduced compared to those generated by the winds of a true cyclone, which has both strong zonal and meridional gradients. However, as the scattered wave energy scales linearly with the incident waves to first order (see (5)), the comparison between the incident and scattered wave-fields remains as an important metric, even if the energy fluxes themselves are under-represented.
@@
Employing the experiment’s zonal and meridional symmetries to separate incident and scattered waves, we calculate the wave energy fluxes propagating outwards through a square centred on $x_{0}$, $y_{0}$ with side lengths $D$. The surface-generated NIW energy flux, $E_{S}$, corresponds to the energy flux propagating meridionally through the northern and southern sides, whereas the energy flux of the NIWs resulting from interactions with topography, $E_{T}$, correspond to the wave energy fluxes through the eastern and western boundaries of the square. Energy fluxes are the product of the velocities $u$ or $v$ normal to the grid-cell face, and pressure $P$. The symmetry of the domain ensures that the energy of the topographically-generated near-inertial waves that are the focus of this study is captured within $E_{T}$.
@@
To isolate the velocities $u$ and pressures $P$ required to calculate $E_{T}$, we subtract reference values from $2000\;\mathrm{km}$ west of the ridge, such that $u^{\prime}=u-u_{\rm far}$ and likewise for $P$. Multiplying the perturbation velocity and pressure, and summing over the opposing sides of the square yields the zonal $E_{i,T}$ and meridional $E_{i,S}$ energy fluxes at the $i$th layer:
$E_{i,T}(y,t)$ $=\left[P_{i}^{\prime}u_{i}^{\prime}\right]_{x=\frac{D}{2}}\;-\;\left[P_{i}^{\prime}u_{i}^{\prime}\right]_{x=-\frac{D}{2}},$ (7)
$E_{i,S}(x,t)$ $=\left[P_{i}v_{i}\right]_{y=\frac{D}{2}}\;-\;\left[P_{i}v_{i}\right]_{y=-\frac{D}{2}}.$ (8)
These energy fluxes are then integrated in space and time to determine the total energy leaving the square meridionally and zonally. The quantities of interest are the vertical energy flux from the storm directly into the surface-generated NIWs, and the vertical energy flux emanating upwards from the topography. Rather than directly calculating these surface and topographic vertical energy fluxes, which is challenging in an isopycnal model with layers intersecting with the topography, we instead assume that these energy fluxes are equivalent to the corresponding horizontal energy fluxes. This assumption means that the calculated energy fluxes will be underestimated by any dissipation or breaking that occurs during the horizontal propagation, but in this idealised model with no background flow and otherwise smooth topography, dissipative mechanisms are minimal. To allow time for the signal to propagate to the edges of the square, the analysis period extends $90$ hours beyond the end of the storm but is still normalised by the $10$ hour storm duration.
With this assumption, we then define our average vertical energy flux quantities $\overline{E_{T}}$ and $\overline{E_{S}}$:
@@ 4 Results
We first provide a qualitative demonstration of the proposed rapid scatter of near-inertial waves at topography. Fig. 3 shows the meridional velocities and isopycnal height anomalies associated with the meridionally propagating near-inertial waves generated directly at the surface. Internal waves are immediately visible through the full water column, similar to the analytic theory shown in Fig. 1. As time progresses, we see that the mode 1 waves (characterised by the single zero crossing) dominant in panel (a) give way to third and fourth vertical modes in panels (c) and (d). This behaviour is consistent with the modal dephasing theory in that the faster, lower modes appear first, which propagate more rapidly from the centre of the storm. By inspecting the phase shifts of individual vertical modes and the beam-like structures, it is clear that the frequency is near-inertial. Importantly to the generation of waves from topography, Fig. 3 shows significant motions at the seafloor immediately after the storm event.
These oscillatory motions are more clearly visible in Fig. 4a, which shows bottom velocities directly beneath the wind forcing similar to the 2D analytic theory in Fig. 1b. Here, the model shows an immediate bottom response with a clear super-inertial signal in the bottom flow, which enables the generation of propagating internal waves in the presence of appropriate topography. A Hovmöller diagram in Fig. 4b shows a signal in the resulting zonal velocity propagating away from the ridge, including several superimposed vertical modes characterised by their different group speeds, fanning out from the topography.
Figure 4: Panel a) Bottom zonal velocity 1000 km West of ridge, b) Hovmöller diagram of zonal velocity anomalies 5 layers from the bottom in the vicinity of the ridge with the approximate group speeds for vertical modes 1, 2, 3, 4, 6, and 9. Both panels show velocities 200 km north of the storm centre, as in the bottom row of Fig. 5.
To visualise the bottom scattered waves, the velocity and interfacial height anomalies along two zonal transects are plotted in Fig. 5. Here, as in (7), the anomalies are calculated with respect to their far-field values to remove the effects of non-topographically-generated waves. In all panels, clear beams can be seen propagating upwards and outwards from the top of the ridge. The two rows, corresponding to locations $150$ and $200\;\mathrm{km}$ northwards from the centre of the wind forcing, show meridional variations to the waves structure and phase. This is due to the time taken for the dephasing signal to propagate outwards from the storm, as well as the meridional structure of the storm itself.
@@
To assess the role of these bottom scattered waves in redistributing energy to higher vertical modes and smaller scales, we take the power spectra of the velocities associated with the meridional and zonal waves in Fig. 6. The prominent vertical modes of the surface-generated, meridional waves visible in Fig. 6a are consistent with the dispersion relation (2), and show a dominant first vertical mode, with energy also spread through modes 2-5. The spectrum of the scattered waves, shown in Fig. 6b, shows bands of power close to harmonics of $|f|$, with local maxima where these bands intersect vertical modes. Compared to the incident meridionally propagating waves, the scattered, zonally propagating waves have a very weak mode 1, and the circled regions in panel b show significant energy present in modes 6-10. Thus energy is distributed further towards higher modes in the scattered zonal waves than the incident meridionally propagating waves. This difference in the modal energy distribution shows that the bottom scatter of NIWs generated at the surface by storms can play a role in downscaling the energy to higher vertical modes, and in turn smaller horizontal scales.
Figure 6: Depth averaged power spectra with the first 10 analytic vertical near-inertial modes overlaid in dashed grey of: a) the meridional velocity associated with the incident surface-generated near inertial waves b) the zonal velocity associated with the topographically scattered waves. Both spectra are taken over the first 10 inertial periods of the experiment, and at the same spatial positions as the corresponding waves shown in figures 3 and 5. The ovals in panel b highlights additional energy in modes 6-10 not present in panel a.
Figure 6a shows that the incident meridional waves have most energy close to $|f|$, but with significant energy in the first harmonic. This enhanced first harmonic can be explained by the wind time scale of the storm – at $10$ hours, the dominant frequency is close to $2|f|$, a choice to ensure that most of the wind energy spectrum is super-inertial, in order to generate the most wave activity. In contrast to Fig. 6a, in Fig. 6b, we see that the first harmonic is more energetic than the near-inertial band, suggesting that this dephasing mechanism for bottom scatter is more efficient for $2|f|$.
This observation is consistent with the strength of the superinertial component of the bottom oscillations compared to the purely inertial in both Fig. 4a and Fig. 1. Here, higher harmonics of $|f|$ are identifiable as the additional, consistently spaced zero crossings between inertial periods. These higher harmonics in Fig. 4a are modulated by by a near-inertial signal of comparable amplitude, much like the two smaller wind forcing scale cases in Fig. 1. The excitation of waves at $2|f|$ is likely due to higher harmonics dephasing more readily than lower: since the spatial scales imposed by the topography remain constant between the $|f|$ and $2|f|$ waves, their group speeds remain the same, resulting in the faster dephasing of the $2|f|$ harmonic due to it having more cycles per unit of distance. This explains the prominence of the $2|f|$ harmonic in this case of a transient storm system generating a rapid bottom response.
The waves shown in Figs. 5 and 4 have characteristics consistent with the proposed mechanism, and Fig. 6 demonstrates their potential for the downscaling of wave energy. Thus, these results serve as a qualitative demonstration of wind-induced topographically generated near-inertial waves. Having qualitatively demonstrated the rapid scatter of storm-generated NIWs at topography, we now quantify the wave energy, in order to assess in what circumstances, if any, these waves might pose a significant source of bottom intensified wave activity.
@@ 4.1 Wave Energy Flux
Having established the mechanism for the rapid topographic scattering of storm generated near-inertial waves in our configuration, we now proceed to calculate the corresponding energy fluxes of the incident and scattered waves. From (7), the time averaged energy flux in our reference configuration (Figs. 3, 5, and 4 above) associated with the wind-generated topographic NIWs is $0.13\;\mathrm{mW}\,\mathrm{m}^{-2}$, compared to $1.9\;\mathrm{mW}\,\mathrm{m^{-2}}$ for the wind-generated surface NIWs. As discussed in section 3, these energy flux values are an underestimate of what might be expected in the real ocean due to the simplified, zonally symmetric wind forcing that enables our clean wave separation. Thus, a more important metric is in comparing the relative sizes of the incident and scattered wave energy fluxes. With the control set of parameters – namely with a ridge $1000\mathrm{m}$ high and $10\mathrm{km}$ wide, and $63~{}\mathrm{ms}^{-1}$ winds lasting for $10$ hours – the zonally propagating, scattered waves have a time-averaged energy flux 7% the size of that of meridionally propagating surface-generated NIWs.
@@ 4.2 Perturbation experiments
We run experiments varying both the properties of the storm and ridge to assess which parameters the energy flux is most sensitive to and determine the conditions under which the topographically-generated waves are most significant.
Figure 7: The energy flux of the wind-driven topographic near-inertial (zonal) waves when varying a) forcing duration; b) surface forcing wind speed; c) ridge height; d) ridge width. The default values are shown with an $\times$. The energy flux for the wind-driven surface-generated near-inertial (meridional) waves is shown in panels a) and b) only, as it is unaffected by varying the shape of the topography.
When altering the duration of the wind forcing pulse (Fig. 7a) there is a peak in both topographic and surface energy fluxes when the forcing duration is less than the inertial period, i.e., $2\pi/|f|\approx 17.5$ hours. This result is consistent with expectations for surface-driven near-inertial waves, as internal wave generation is only supported by super-inertial forcing frequencies. Forcing durations right at the inertial frequency would overlap significantly with the sub-inertial band, whereas a more rapid surface forcing would contain less energy at the highly resonant inertial period. The offset peaks in Fig. 7a) are likely due to the exciting of the $2|f|$ harmonic in the scattered waves, shown in Fig. 6b. This is due to the significant $2|f|$ component of the bottom velocity (see Fig. 4a) which in turn is caused by the faster dephasing of the higher harmonic waves than the near-inertial.
@@
Finally, Fig. 7d shows the sensitivity of topographic energy flux when varying the ridge width. Here, the scaled energy flux, as in (7), is shown alongside the product of the energy flux and width. This plot is designed to show the way that the total wave energy changes, as well as energy flux, which scales inversely to ridge width. Here, we see that the energy flux per unit area (Fig. 7d, left axis) is greatest for the narrowest ridge, but that the total wave energy per metre meridionally (Fig. 7d, right axis) peaks for a $30\;\mathrm{km}$ wide ridge. This effect is due to the maximum topographic slope becoming shallower than the angle of the wave beam, suppressing the downward propagating part of the wave-field visible in Fig. 5 (result not shown). In (5), $k$ refers to the dominant wavenumber of the ridge, which in our case is $k\sim 1/W$. This is consistent with the our results in Fig. 7, whereby the energy flux per unit area is greater for steeper and narrower topography.
@@ 5 Discussion and conclusions
We demonstrate a novel mechanism for the generation of wind-driven near-inertial waves at topography within hours of the passage of a storm. These waves are generated by the rapid dephasing of horizontally propagating full-depth wave modes induced by a transient wind event, and interaction of the resulting bottom flow with topography to radiate waves. This generation mechanism is showcased using an isopycnal model forced by a wind pulse that mimics a storm. The idealised model setup is carefully constructed to enable a clean separation of the scattered waves from surface-generated near-inertial waves and to eliminate any significant time-mean flow that could generate lee waves.
@@
The perturbation experiments shown in Fig. 7 reveal that the variation of wave energy fluxes in our experiment is consistent with the existing internal wave theory. With our control experiment, namely a $1000\;\mathrm{m}$-high, $10\;\mathrm{km}$ wide ridge in a $4000\;\mathrm{m}$ deep ocean, with an $80\;\mathrm{km}$ wide (FWHM) category 4 tropical cyclone overhead, the topographically scattered near-inertial waves have energy fluxes of $0.13\;\mathrm{mW}/\mathrm{m}^{2}$, that is, $7\%$ the energy flux in surface-generated NIWs in the same experiment. Given that our simplified storm under-represents surface NIW generation, we focus on the comparison between the surface-generated and topographically scattered waves, and consider these energy flux estimates as a lower bound.
With the most narrow topography that we test, the maximum energy flux attained is $0.39\;\mathrm{mW}/\mathrm{m}^{2}$, or $18\%$ of the surface-generated NIWs. However, as such a prominent ridge or seamount would be uncommon, we consider as a more typical range, the energy flux ratios from the strength and duration perturbation experiments (Fig. 7a,b). From these two experiments, the ratios of the topographically generated to the directly surface-generated NIW energy fluxes range from 5% at the longest duration and strongest storm, to 10% at the shortest and weakest storms. This range is consistent with the 6% of surface NIW energy estimated to reach the seafloor in the work by Jouanno et al. (2016), except that with our proposed mechanism this energy is communicated rapidly (within 100 hours of the storm) to the bottom.
@@
Given that the mechanism described in this study relies on specific and short-lived circumstances, namely a storm event near prominent topography, it is not expected to be a major contributor to the internal wave field in a globally averaged sense. Instead, the mechanism could be important for explaining intermittent locally enhanced internal wave generation at the seafloor and deep-ocean mixing. Recent studies have identified evidence for unexplained enhancement of internal wave activity at the seafloor. van Haren (2020) finds anomalously large bottom near-inertial oscillations, mixing and deep overturning at $3000\;\mathrm{m}$ depth $1000\;\mathrm{km}$ away from a category 4 tropical cyclone. This bottom signal’s first detection and intensification happened simultaneously with the approach of the cyclone, pointing to a rapid (on the order of hours or a few days) communication of the storm’s energy to the deep ocean. Similarly, in another observational study, Morozov and Velarde (2008) show a rapid bottom response coinciding with the arrival of two cyclones, in addition to the downward propagating wave-packet which takes two weeks to reach the seafloor (see Figs. 11 and 12 by Morozov and Velarde (2008)). A modelling study of storm-induced flows shows enhanced near-inertial energies at the seafloor within 5 days of the passage of a storm, several days ahead of the downward propagating wave packets (see Fig. 9 by Jouanno et al. (2016)). Jouanno et al. (2016) note that while this bottom NIW generation is linked to the storm, the associated energies are two orders of magnitude smaller than the surface-generated NIWs and were thus not investigated further. However, our perturbation experiments suggest this ratio of surface to bottom generated NIW energy flux could be an order of magnitude larger than was found by Jouanno et al. (2016), warranting further investigation. We argue that the mechanism proposed in the study could at least partly account for these findings in the existing literature, but further study is needed to clarify the extent to which this may be the case.
Here, we studied the generation of wind-driven topographic near-inertial waves in a highly idealised domain to cleanly separate them from other flow features. An important next step is to identify the waves in more realistic simulations. However, delineating topographic near-inertial waves from other flow factors is difficult due to their generation by the dephasing of the vertical modes of the surface response, which itself includes NIWs of similar frequencies and spatial scales. Our experiments were carefully designed to separate the wind-driven topographic near-inertial waves from other flow features as cleanly as possible. Consequently, this caused the underestimation of the near-inertial wave energy imparted by the storm and limited the scope of our study to simple dynamics, topographic features, and surface forcing profiles. To investigate topographic near-inertial waves in a more realistic numerical simulation or with observations, these intricacies of disentangling them from other flow features will need to be handled carefully. One option could be to exploit the difference in vertical propagation between the topographic and wind-generated near-inertial waves. If a filter were used to separate the upwards and downwards propagating signals, perhaps following the spectral approach employed by Olbers and Eden (2017) or Waterhouse et al. (2022), one could identify areas of increased vertical energy flux beneath storm systems. However, one would still need to remove the background internal wave fields, among other oscillatory signals that could obscure the waves of interest. To obtain a stronger signal, a statistical approach could be employed, e.g., averaging over many storm events in several locations above regions of varying topographic prominence. At each location, measurements of the bottom mixing could be taken both in the absence and presence of an overhead storm. Future studies could employ this approach on high-resolution model output or observations.
As shown in Fig. 6, the scatter of storm-generated NIWs at topography can redistribute energy towards higher vertical modes, and smaller horizontal scales. The higher vertical modes visible in Fig. 6b suggests that these scattered waves may be especially important for abyssal mixing due to their role in energy downscaling. However, our study was designed to maximise both wave generation and propagation so that the scattered wave-field was easier to isolate and study and therefore mechanisms for wave breaking and dissipation were intentionally either removed or suppressed. Hence we made no attempt to estimate abyssal mixing generated by this mechanism.
In addition to the spatial downscaling, the scattered waves tend to higher frequencies, particularly the prominent first harmonic of $|f|$. The reason for this is that the power spectrum of the bottom velocity from the dephasing of the incident waves in Fig. 4 contains a $20\%$ stronger $2|f|$ peak than $|f|$ (figure not shown), likely because the higher harmonics dephase more quickly and so are more prominent in the post-storm bottom velocity. As higher frequency internal waves exhibit less vertical shear, they are less prone to breaking, so this could somewhat offset any increase to mixing attributed to the spatial downscaling. Higher harmonics in the downscaling of near-inertial waves by eddies was observed in Danioux et al. (2011). In this study, authors found that the $2|f|$ harmonic dominated $|f|$ at depth, stating that the mechanism explaining this shift in frequency in depth remains unresolved. If the role of eddies in downscaling these near-inertial waves also involves the dephasing of vertical modes, then the results of our study could help to explain these findings.